OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = 1/3x - 3.

OpenStudy (anonymous):

first correct answer gets a medal

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

y = 3x - 14 y = 1/3x + 2 y = -3x + 22 y = - x + 8

OpenStudy (anonymous):

thank you alot